part c and d pleaseand y= 1-1 + i 4 5-i If x=5i 1-i find the following. (a) xy 18i (b) yły -12i+ 40 (c) (x, y) 18i (d) (y, y) 40 - 12i
Answer #1

15 2- (3) 18-63). 5- cy (28) = 4x (1-1) + 58(4) + (-i)(5+1) =-4-41+ zoits-si tiri = -4 -4 201 +5 - diti - 2+12 1 dy lý, %) =